### v1

# Zero-shot Context Biasing with Trie-based Decoding using Synthetic Multi-Pronunciation

Abstract
Contextual automatic speech recognition (ASR) systems allow for recognizing out-of-vocabulary (OOV) words, such as named entities or rare words. However, it remains challenging due to limited training data and ambiguous or inconsistent pronunciations. In this paper, we propose a synthesis-driven multi-pronunciation contextual biasing method that performs zero-shot contextual ASR on a pretrained Whisper model. Specifically, we leverage text-to-speech (TTS) systems to synthesize diverse speech samples containing each target rare word, and then use the pretrained Whisper model to extract multiple predicted pronunciation variants. These variant token sequences are compiled into a prefix-trie, which assigns rewards to beam hypotheses in a shallow-fusion manner during beam-search decoding. After which, any recognized variant is mapped back to the original rare word in the final transcription. The evaluation results on the Librispeech dataset show that our method reduces biased word error rate (WER) by 42% on test-clean and 43% on test-other while maintaining unbiased WER essentially unchanged.
Index Terms: ASR, TTS, Contextual biasing, Zero-shot.
1 Introduction
Contextual automatic speech recognition (ASR), also known as “hotword” or ”named-entity” customization, enables ASR systems to accurately recognize user-specified terms, such as names, places, or technical jargon, that are rare or even out-of-vocabulary (OOV) in training data [1]. Modern end-to-end (E2E) ASR models achieve an impressive overall word error rate (WER) compared to traditional hybrid systems, thanks to advanced architectures such as RNNs [2], Transformers [3], and Conformers [4]. However, due to data scarcity and confusion in pronunciation for rare words, contextual ASR remains challenging [5].
Traditional contextual biasing techniques augment decoding via shallow‐fusion with an external language model [6] or on-the-fly WFST rescoring [7]. While effective, these methods require careful tuning of fusion weights and can degrade WER/U-WER performance when the bias list becomes large, due to an increase in false positive distractor terms with similar spellings or pronunciations. To be more straightforward, the E2E approaches embed context phrases directly into the ASR model, for example, Contextual Listen, Attend and Spell (CLAS) uses an attention‐based contextual module to bias decoding [8], and trie-based deep biasing integrates dynamic n-gram constraints into transducer decoding [9]. Recently, large language model (LLM)-based ASR has emerged, leveraging prompt engineering to incorporate hotwords into decoding. For example, a CTC-assisted LLM-based contextual ASR model uses coarse CTC hypotheses to select effective hotwords from a large hotword list, which are then fed into an LLM as a prompt [10]. These yield substantial improvements in biased-WER (B-WER) by enhancing the model’s capability for contextual understanding during the inference phase.
Despite these advances, one critical factor has often been overlooked: the variability in pronunciation of hotwords. In practice, many hotwords, especially specialized terms such as technical jargon, foreign names, or branded entities, are unfamiliar to users, leading to inconsistent or incorrect pronunciations. As a result, the same hotword may be spoken in multiple ways across different speakers or even by the same speaker over time. This mismatch between the intended hotword and its spoken realization can significantly degrade the effectiveness of contextual biasing, as models may fail to associate these variant forms with the intended target.
In this paper, we introduce a novel approach to enhancing the recognition accuracy of rare words in a pretrained ASR system without requiring any adaptation. Our method leverages synthetic speech to generate diverse pronunciation variants for each hotword, which are then transcribed by the same ASR model. Technically, we synthesize speech using a variety of TTS systems and voices for each templated hotword sample and apply Whisper-large-v3 model to transcribe them. We integrate a multi-pronunciation prefix-trie-based hotword module into the same Whisper-large-v3 model in a shallow-fusion manner, where the prefix-trie is constructed using extracted multiple pronunciation variants of each hotword.
Our experiments on the Librispeech test sets demonstrate that the proposed method reduces WER/B-WER of the baseline Whisper-large-v3 model without affecting the recognition performance of common words. To the best of our knowledge, this is the first work to leverage synthesized multi-pronunciation variants for contextual ASR without requiring model adaptation.
2 Methodology
In this section, we describe our end‐to‐end pipeline for augmenting Whisper’s contextual biasing with TTS‐generated multi-pronunciation variants. The pipeline comprises three sequential stages: Hotword Speech Synthesis, Hotword Token Extraction, and Prefix-Trie Implementation in Whisper.
Figure 1: Workflow of TTS-powered multi-pronunciation prefix-trie construction for Whisper model. Given a list of hotwords, each is inserted into a designated template and synthesized using multiple TTS engines to mimic three distinct speaker voices. The synthetic speech is then transcribed using the vanilla Whisper model, from which diverse pronunciation variants are extracted. Each unique transcription is included as an effective hotword in the prefix-trie. During decoding, any recognized variant shown in the final results is mapped back to the original hotword.
2.1 Multi-Pronunciation Hotword Speech Synthesis via TTS
We begin by synthesizing multiple pronunciation variants for each target hotword using three complementary TTS engines, including CosyVoice [11], F5-TTS [12], and GPT-SoVITS [13].
  • F5-TTS is a non-autoregressive (NAR) TTS model that builds on a Diffusion Transformer (DiT) architecture and incorporates ConvNeXt V2 blocks to effectively address text-speech alignment in an in-context learning setting. It is trained on 95K hours of English-Chinese bilingual dataset and supports good-quality synthesis in both languages.
  • CosyVoice is a scalable zero-shot multilingual TTS system that employs a text-to-token-to-speech architecture. It leverages supervised semantic tokens extracted from a quantized multilingual ASR encoder, uses a large language model (LLM) to predict token sequences from text and speaker embeddings, and reconstructs speech via a conditional flow matching model followed by a HiFi-GAN vocoder. It is trained on around 170K hours of multilingual dataset and supports five languages and dialects, including Chinese, English, Cantonese, Japanese, and Korean.
  • GPT-SoVITS integrates a speech-to-unit encoder, a speaker encoder, and a GPT-style text-to-unit decoder within the So-VITS framework. It enables prompt-based zero-shot voice cloning by generating discrete unit sequences from text and synthesizing them into a waveform via a neural vocoder. It supports the same five languages as Cosyvoice.
All three systems support prompt-based speech synthesis by conditioning on a reference audio, its corresponding text, and a new target text, allowing for zero-shot speech generation in the reference speaker’s voice and style.
To ensure that the hotword appears in a fixed, easily locatable context, we embed it in two minimal utterances: “Start <Hotword> End” and “Begin <Hotword>”. The context can be customized, but we adopt such simple and rigid templates to facilitate reliable extraction from Whisper decoding results and to ensure stable and consistent synthesis from TTS models. Longer or more complex contexts tend to introduce unwanted variability in TTS outputs, which can compromise the reliability of anchor-based hotword localization, as the surrounding context serves as a temporal marker for extracting hotword speech variations. Each engine renders both utterances with three distinct voices and accents from [14], including female British accent, male British accent, and male American accent. This process yields a total of 3 engines × 2 utterances × 3 voices = 18 synthesized utterances per hotword, capturing a rich and comprehensive variety of pronunciation patterns.
2.2 Multi-Pronunciation Hotword Token Extraction
Figure 2: An illustration of Single-Pron-Hotword and Multi-Pron-Hotword prefix-trie using Final-token-only and Uniform per-token reward schemes. In Single-Pron-Hotword prefix-trie, each hotword corresponds to only one path, standing for the correct way of spelling. In contrast, in a Multi-Pron-Hotword prefix-trie, each hotword may correspond to multiple paths with different spellings. When using the Final reward, only the last state of a valid path will give a reward of 1; whereas when using Uniform reward, each active state will credit a reward of 1 when matched. Numbers in the parentheses stand for the rewards gained till the current state. Red dashed lines are examples of failing to match in the middle state of a prefix-trie path.
Next, each synthesized waveform is fed into the vanilla Whisper-large-v3 model to obtain the recognition results, where we require not only the final sentences but also the sequences of byte-pair encoding (BPE) tokens. This is to align with the requirement of prefix-trie as mentioned in Section 2.3, which performs hotword biasing through BPE-level rewarding scores in each auto-regressive decoding step of the Whisper decoder that generates the probability distribution of the entire BPE dictionary.
From the token sequences, we locate the tokens corresponding to the known context markers “Start,” “End,” and “Begin,” and we isolate the subsequences that occur between “Start” and “End” as well as those immediately following “Begin”. Any extracted sequence whose token IDs exactly match the original hotword is discarded, so that only unmatched pronunciation-prediction variant pairs remain. This extends the search space to include more predictions that the Whisper model can make for each hotword, allowing us to improve hit accuracy. Additionally, we refine the variant set to eliminate faulty options by Syllable Filtering (SF). Specifically, we use the syllapy [15] package to estimate the number of phonemes of the predicted variants. Only those with syllable count matches that of the original hotword and at least three are retained. This strategy significantly improves performance by removing inappropriate pronunciation variants introduced by TTS generation instability or the Whisper model’s hallucinations through the automated pipeline.
2.3 Prefix-Trie Implementation in Whisper Decoder
We implement a simple prefix-trie module into the Whisper beam-search decoding to enable contextual biasing in Whisper without requiring model adaptation or other learnable modules, such as Tree-constrained Pointer Generator (TCPGen) [16]. In our prefix-trie architecture, each arc, corresponding to a token, is annotated with a Uniform per‐token reward of +1+1, so that every token in every pronunciation variant contributes equally to decoding. Since each token in a hotword contributes a fixed reward during decoding, longer hotwords (i.e., those with more tokens) inherently receive more cumulative bias, giving them a greater advantage when matched correctly. During beam‐search decoding, we maintain, for each beam hypothesis, a corresponding “active” trie state that records how far along any hotword prefix that hypothesis has traversed. Given the audio $X$, at each time step $t$, when considering an output token $y_{t}$, we first compute the Whisper score $S_{\textbf{W}}$ as in Eq. 1.
${S_{\textbf{W}}}(y_{t})=-\log P_{\textbf{W}}(y_{t}\mid y_{t-1},...,y_{0},X)$ (1)
Here we define the cumulative score of each beam Hypothesis as in Eq. 2.
${S_{\textbf{H}}}(y_{t})={S_{\textbf{H}}}(y_{t-1})+{S_{\textbf{W}}}(y_{t})$ (2)
We then attempt to advance the hypothesis’ current trie state with $y_{t}$: if $y_{t}$ continues a valid hotword prefix conditioned on $y_{t-1}$, the trie returns a reward $\rho_{y_{t-1}\rightarrow y_{t}}=1$; otherwise we remove the accumulated reward of partial unsuccessful match from the hypothesis’s cumulative score. We define the reward as $\sum_{i=1}^{n}\rho_{y_{t-i}\rightarrow y_{t-i+1}}$ where $n$ is the depth from root to current state of prefix-trie as shown in Figure 2. We form the final score $S$ augmented with contextual rewards, as in Eq. 3, which is then used in place of $S_{\textbf{W}}$} when ranking and pruning beam candidates.
$S(y_{t})={S_{\textbf{W}}}(y_{t})+\rho_{y_{t-1}\rightarrow y_{t}}$ (3)
In this way, any hypothesis that follows one of our multi‐pronunciation hotword paths receives a consistent positive boost, biasing the search toward recognizing those variants. After the beam-search decoding, we select the highest‐scoring beam, and then traverse its recorded trie states to map to the correct hotword spelling. Since we only add these Uniform per-token rewards to the beam scores without altering the underlying acoustic or language model logits, our trie‐based biasing improves recognition of the target hotwords while preserving Whisper’s overall decoding behavior and general ASR accuracy.
Through this three-stage methodology, as shown in Figure 1, we incorporate a prefix-trie-based multiple pronunciation hotword mechanism with Whisper for contextual biasing, leading to significant reductions in biased-WER without degrading overall ASR accuracy.
3 Experiments
In this section, we introduce the dataset and evaluation metrics used, as well as the experimental setup.
3.1 Dataset and Evaluation Metrics
3.1.1 Data and Hotword Selection
We follow previous research [9] on their settings regarding contextual biasing and benchmark our method on the widely used Librispeech dataset [17], including the test-clean and test-other evaluation sets. Specifically, we adopt the artificial biasing word list setup [^1] The rare and common word lists are from https://github.com/facebookresearch/fbai-speech/tree/main/is21_deep_bias. where the 5,000 most frequent words from the Librispeech training data are labeled as common, while all remaining words are treated as rare, including over 209.2K words. For each test utterance, the biasing list is composed of two components: rare words that appear in the reference transcription, and distractor words randomly selected from the rare words list. We use a fixed $N=1000$ distractors to align with prior works for a fair comparison. The statistics of both test sets are as demonstrated in Table 1. Rare words take up 10% of the word count in both test sets.
From the rare word vocabulary, we select a subset of rare words that exhibit a non-zero initial word-error rate (WER) when decoded with the official Whisper-large-v3 model without contextual biasing. We then perform speech synthesis on this subset. In contrast, rare words that are correctly transcribed are explicitly excluded from further processing and do not go through the synthesis pipeline. Both sections of rare word vocabulary above are taken into account when calculating biased-WER.
Table 1: Statistics of the Librispeech dataset. Appearing rare words refers to the rare words counts that are present in the reference transcripts.
Metric test-clean test-other
Total words 52,576 52,343
Common words 46,815 46,993
Appearing rare words 5,761 5,350
Utterances 2,620 2,930
Avg. rare words per utterance 2.20 1.83
Rare word rates 10.96% 10.22%
3.1.2 Evaluation Metrics
We utilize commonly used evaluation metrics for contextual ASR tasks: overall WER to measure general recognition quality, biased-WER (B-WER) to specifically evaluate the accuracy on targeted hotwords, and unbiased-WER (U-WER) to measure the accuracy on common words. Comparing contextual ASR with standard ASR systems, the combination of these three WERs shows the effectiveness (true positives) and distractions (false positives) of contextual biasing strategies embedded in the ASR systems. Thus, an optimal contextual biasing ASR should achieve improved B-WER while maintaining U-WER unchanged. For transcription normalization and metrics calculation, we utilize the scripts provided by MetaAI [9].
3.2 Experimental Setup
All of our experiments are based on the Whisper-large-v3 model, and we use beam-search decoding with a beam size of 10. We use the vanilla Whisper-large-v3 model as a baseline, comparing against the Whisper-large-v3 with prefix-trie hotword module under various configurations as illustrated in Figure 2:
  • Single-Pron-Hotword: Each hotword is bound with only one way of spelling, which is the original correct one, and corresponds to only one path in the prefix-trie.
  • Multi-Pron-Hotword: As proposed in Chapter 2, each hotword is bound with several selected pronunciations generated through a TTS-Whisper pipeline, and may correspond to several paths in the prefix-trie.
Furthermore, we explore two reward strategies in the prefix-trie: a Final-token-only reward scheme that gives a reward of $1$ only when the decoding beam reaches the terminal node of the hotword in the trie, and a Uniform per-token reward scheme where each hit token of the hotword returns a reward of $1$. The Final scheme only rewards those beams where Whisper predicts the entire correct sequence of hotwords on its own. In contrast, the Uniform scheme continuously rewards a beam each time a partial token of a hotword is predicted; however, if at any point the next expected token does not match, the awarded reward is immediately removed.
4 Results
Table 2: Performance of proposed multi-pronunciation hotword on Librispeech dataset using biasing list with $N=1000$. SF indicates whether we apply syllable filtering to the multi-pronunciation variants. Reward indicates the reward patterns (Final-token-only vs. Uniform per-token) used. Final-token-only indicates only the final prefix-trie state of a valid path has a reward of 1 while the previous states has a reward of 0. Uniform per-token indicates every prefix-trie state of a valid path has a reward of 1. WER refers to the overall word error rate. B-WER/U-WER refers to biasing/unbiasing WER, indicating the model’s ability to recognize rare/common words.
System Condition SF Reward test-clean test-other
WER U-WER B-WER WER U-WER B-WER
Baseline Vanilla Whisper-large-v3 - 2.81 1.94 9.86 4.84 3.39 17.55
S1 Single-Pron-Hotword Final 2.57 1.91 7.95 4.27 3.21 13.61
S2 Multi-Pron-Hotword Final 19.97 21.19 9.98 26.45 27.25 19.42
S3       + Single-Pron Final 19.79 21.17 8.54 26.15 27.21 16.84
S4 Multi-Pron-Hotword Final 2.70 1.96 8.73 4.52 3.25 15.68
S5       + Single-Pron Final 2.50 1.92 7.19 4.18 3.22 12.60
S6 Single-Pron-Hotword - Uniform 2.33 1.87 6.11 3.87 3.11 10.52
S7       + Multi-Pron Uniform 2.31 1.90 5.66 3.83 3.14 9.90
Table 3: Comparison with prior contextual ASR models on Librispeech dataset using biasing list [9] with list size N=1000.
Model test-clean test-other
WER U-WER B-WER WER U-WER B-WER
CPPNet [18] 3.81 2.90 11.40 8.75 6.90 25.30
Deep Biasing+BPB [19] 3.47 3.00 7.70 7.34 6.40 15.80
Whisper+TCPGen+GPT-2 [20] 3.40 8.20 6.30 16.30
TCPGen+GNN enc. [21] 3.10 6.70 7.90 17.80
GA-CTC [22] 2.40 2.00 6.30 6.20 5.20 15.20
TCPGen+p+phn-aware Q [23] 2.20 4.60 6.00 12.30
DB-NNLM [9] 2.14 1.60 6.70 6.35 5.10 17.20
CTC+LLM [10] 1.33 1.00 4.16 2.99 2.31 9.33
S7 2.31 1.90 5.66 3.83 3.14 9.90
Table 2 summarizes our evaluation results on Librispeech test-clean and test-other subsets. Firstly, we test the Final reward scheme to evaluate Whisper’s ability to recognize rare words, as we only reward hypotheses that contain correctly recognized hotwords and prioritize these decoding paths. Then, SF is applied to improve the reliability of the multi-pronunciation hotword list. Finally, we utilize the Uniform reward scheme to guide Whisper’s decoding process toward hotword recognition by providing rewards for partial matches, thereby introducing a bias in favor of contextual biasing.
In Table 2, with the reward scheme set to Final, comparing S1 and the Baseline system, we observe a significant B-WER reduction (19% for test-clean and 22% for test-other), along with slightly better U-WER results. This suggests that a prefix-trie hotword implementation already provides substantial improvements in hotword recognition for Whisper. When we implement the multi-pronunciation hotword strategy without using SF (as in S2 and S3), the U-WER performance degrades dramatically. At the same time, B-WER does not improve at all, showing that our pipeline for generating the multi-pronunciation hotwords brings us a large amount of noisy variations, leading to a wrong rewarding paradigm. However, when SF is applied to remove those faulty variations, comparing S5 with S1, we can conclude that the strategy using multi-pronunciation hotwords with their original forms introduces further B-WER improvement by 10% and 7% for test-clean and test-other, respectively, while the U-WER remains essentially unchanged. Since the rare words only occupy around 10% of the entire test sets, S5 only achieves 2-3% improvement on the overall WER.
Using Uniform reward scheme, comparing S6 with S1, U-WER and B-WER have identical improvement for both test sets, with around 3% performance improvement on U-WER, and 23% B-WER reduction. Also, we observe the same trends when comparing S7 against S5. This indicates that guiding Whisper’s decoding process by rewarding step by step to flavor rare words achieves even better performance. Our final system, S7, integrates a prefix-trie contextual biasing module built on multi-pronunciation and the original form of rare words to the Baseline system, achieving 18%, 21% overall WER reduction, and 43%, 44% B-WER performance improvement, for test-clean and test-other, respectively.
We also compare our method against many prior contextual ASR approaches evaluated on the Librispeech dataset with biasing list from [9] in table 3. Despite utilizing a frozen Whisper backbone without fine-tuning or LLM assistance, our proposed method significantly outperforms several fine-tuned models in terms of B-WER under the same configuration of biasing list and distractor size ($N=1000$).
5 Conclusion
We introduced a novel approach to enhancing zero-shot contextual ASR performance by leveraging synthetic speech for Whisper. Our method involves synthesizing diverse hotword speech via TTS engines and integrating these multi-pronunciation variants extracted from the Whisper model into a prefix-trie for shallow-fusion in beam-search decoding. Without any model fine-tuning or reliance on large language models, our system consistently improves hotword recognition performance on the Librispeech test sets with a hotword biasing list, achieving competitive biased-WER while maintaining and even improving general ASR quality. Our results surpass many fine-tuned contextual ASR systems, making the approach practical, low-cost, and easy to deploy, as it does not require any fine-tuning of the underlying ASR model. Future work will expand the TTS augmentation pipeline by incorporating additional voice styles and more diverse TTS engines, including multilingual and expressive speech synthesis. This would enable coverage of broader pronunciation variability and better simulate real-world user scenarios across regions and accents.